# Unleashing the Potential of Vision-Language Pre-Training for 3D Zero-Shot Lesion Segmentation via Mask-Attribute Alignment

### v1

@@ 4.3 Ablation Studies
Figure 3: Visualization illustrating how utilizing both mask and text embeddings improves segmentation performance for both unseen and seen lesions. We present several prediction cases generated using both text embeddings and mask tokens, text embeddings only (TE), and mask tokens only (MT).
Ablation of the Cross-Modal Knowledge Injection module. We examine key components in the proposed Cross-Modal Knowledge Injection module in detail, including (1) the significance of deep fusion (DF), which is designed to enhance the representation of mask tokens and text embeddings; (2) the effectiveness of leveraging both text embeddings (TE) and mask tokens (MT) to generate predictions. The results are shown in Table 4. Comparing the results of the first three rows with the last three rows (highlighted in light red), it is evident that deep fusion significantly improves performance, whether using only text embeddings, only mask tokens, or both for segmentation result prediction. This observation shows the importance of enabling cross-modal information interaction for optimal performance. Furthermore, whether or not deep fusion is applied, using both text embeddings and mask tokens for segmentation prediction, combined with ensembling the results, consistently outperforms using only unimodal token embeddings for segmentation prediction. This confirms the superiority of leveraging the complementary strengths of both visual and textual embeddings to further enhance segmentation performance. Fig. 3 provides detailed examples illustrating how combining text tokens and mask tokens improves performance. While using mask tokens for segmentation prediction is more intuitive and generally yields better performance than using text embeddings alone (as shown in Table 4), we observed that when lesion boundaries are particularly blurry (as seen in the Hepatic Vessel Tumor cases on the left side of Fig. 3), the visual features from the mask tokens fail to accurately capture the lesion’s area. In such instances, text descriptions of attributes—such as the lesion’s shape and surface characteristics—provide additional information. The segmentation generated by the corresponding text embeddings refines the results produced by mask tokens alone, leading to improved performance. Moreover, we observed that, owing to the clear semantic information and specific contexts carried by the textual features of attribute descriptions, the segmentation results generated using attribute embeddings exhibit significantly fewer false positives. This further refines the segmentation produced by mask tokens by reducing false positives, as demonstrated in the colon tumor cases on the right side of Fig. 3.
@@ Appendix A Preliminaries
Mask Tokens. Recently, a transformer-based segmentation model Mask2Former (Cheng et al., 2022) that can generate a set of segment-level embeddings have been applied and achieved promising performance in the field of medical image segmentation (Yuan et al., 2023; Chen et al., 2023; Jiang et al., 2024). Different from standard segmentation methods, Mask2Former adopt an additional lightweight transformer decoder along with a vision encoder and a pixel decoder. By feeding $N$ mask tokens (also called segment queries) and pixel decoder features into the transformer decoder, Mask2Former generates $N$ segment-level embeddings with masked attention layers, effectively partitioning an image into $N$ corresponding regions. The final binary mask segmentations are obtained by computing the dot product of the mask tokens with the image features from the last pixel decoder layer, while an additional MLP predicts the class for each mask. In contrast, this work extends the closed-set setting of mask classification to an open-set setting through fine-grained mask-text alignment. Specifically, we propose enhancing and matching mask tokens with text embeddings of the visual attributes of pathologies within multi-scale transformer decoder blocks. Additionally, we leverage the complementary strengths of visual and text representations for mask refinement, demonstrating strong generalizability and improved zero-shot lesion segmentation performance.
@@ B.2 Creating Structured Disease Attribute Descriptions
Specifically, For the public datasets MSD and KiTS23, which only provide mask annotations, two senior radiologists supplemented the annotated lesions with corresponding lesion descriptions of the eight visual attribute aspects. For the in-house data, we utilized the semi-automatic annotation pipeline described in Sec.3.1. First, we prompted GPT-4(Achiam et al., 2023) to extract lesion-related descriptions from the findings section of each report. Then, two experienced radiologists collaborated to review, correct, supplement, and expand the GPT-generated visual attribute descriptions. Through this process, we established eight standardized attributes for describing lesions, which form the basis of our structured reports. These attributes provide a comprehensive and objective template for characterizing lesions, ensuring consistency and enabling generalization across different types of lesions, as shown in Table 6.
@@ Appendix C Fine-Tuning Evaluation
Appendix D Evaluation of Lesion-Attribute Matching
@@
Figure 4: Visualization of segmentation and lesion-attribute matching results of Malenia. The numbers indicate the normalized similarity scores between the lesion masks and the corresponding attributes.
Appendix E Computational Cost Comparison
@@
Appendix F Limitations and Future Work
@@
Malenia currently achieves superior zero-shot lesion segmentation performance, approaching that of fully-supervised, task-specific models. However, zero-shot lesion segmentation across anatomical regions and imaging modalities remains highly challenging. For instance, training on lesion data from abdominal CT and directly testing on brain cancer MRI is extremely challenging due to the significant differences in anatomical structures and feature distribution between the training and testing images. We believe that achieving cross-domain and cross-modality zero-shot generalization is a challenging yet meaningful research direction in both natural and medical imaging domains. This will be the focus of our future work.

### v2

@@ 4.3 Ablation Studies
Ablation of the Cross-Modal Knowledge Injection module. We examine key components in the proposed CMKI module in detail, including (1) the significance of deep fusion (DF), which is designed to enhance the representation of mask tokens and text embeddings; (2) the effectiveness of leveraging both text embeddings (TE) and mask tokens (MT) to generate predictions. The results are shown in Table 4. Comparing the results of the first three rows with the last three rows (highlighted in light red), it is evident that deep fusion significantly improves performance, whether using only text embeddings, only mask tokens, or both for segmentation result prediction. This observation shows the importance of enabling cross-modal information interaction. Furthermore, whether or not deep fusion is applied, using both text embeddings and mask tokens for segmentation prediction, combined with ensembling the results, consistently outperforms using only unimodal token embeddings for mask prediction. This confirms the superiority of leveraging the complementary strengths of both visual and textual embeddings to further enhance segmentation performance.
@@ Appendix A Preliminaries
Mask Tokens. Recently, a transformer-based segmentation model Mask2Former (Cheng et al., 2022) that can generate a set of segment-level embeddings have been applied and achieved promising performance in the field of medical image segmentation (Yuan et al., 2023; Chen et al., 2023; Jiang et al., 2024). Different from standard segmentation methods, Mask2Former adopt an additional lightweight transformer decoder along with a vision encoder and a pixel decoder. By feeding $N$ mask tokens (also called segment queries) and pixel decoder features into the transformer decoder, Mask2Former generates $N$ segment-level embeddings with masked attention layers, effectively partitioning an image into $N$ corresponding regions. The final binary mask segmentations are obtained by computing the dot product of the mask tokens with the image features from the last pixel decoder layer, while an additional MLP predicts the class for each mask. In this work, we extend the closed-set setting of mask classification to an open-set setting through fine-grained mask-text alignment. Specifically, we propose enhancing and matching mask tokens with text embeddings of the visual attributes of pathologies within multi-scale transformer decoder blocks. Additionally, we leverage the complementary strengths of visual and text representations for mask refinement, demonstrating strong generalizability and improved zero-shot lesion segmentation performance.
@@ B.2 Creating Structured Disease Attribute Descriptions
Specifically, For the public datasets MSD and KiTS23, which only provide mask annotations, two senior radiologists supplemented the annotated lesions with diagnostic reports and corresponding lesion descriptions of the eight visual attribute aspects. For the in-house data, we utilized the semi-automatic annotation pipeline described in Sec.3.1. First, we prompted GPT-4 (Achiam et al., 2023) to extract lesion-related descriptions from the findings section of each report. Then, two experienced radiologists collaborated to review, correct, supplement, and expand the GPT-generated visual attribute descriptions. Through this process, we established eight standardized attributes for describing lesions, which form the basis of our structured reports. These attributes provide a comprehensive and objective template for characterizing lesions, ensuring consistency and enabling generalization across different types of lesions, as shown in Table 6.
@@ B.3 Data Preprocessing
B.4 Justification for the Legality of Data Use
This study utilizes data from both public and private sources. (1) The public data is obtained from the MSD (Antonelli et al., 2022) and KiTS23 (Heller et al., 2023) datasets, which are publicly available and provide CT images along with lesion masks. We have supplemented these lesion masks with corresponding structured lesion attribute descriptions. Since these datasets are fully open and no patient privacy information was disclosed during their use, our data usage is fully justifiable and legal. (2) The private dataset is sourced from the First Hospital of China Medical University. This retrospective study was approved by the Ethics Committee of the First Hospital of China Medical University, and the informed consent requirement was waived. The private dataset includes CT images, lesion masks, and diagnostic reports. All diagnostic reports in this dataset have been fully anonymized, with all patient-identifiable information and private details removed. For all anonymized diagnostic reports, we exclusively utilized the GPT-4 API (Achiam et al., 2023) to process textual descriptions of lesions, converting them into a structured format. No patient-identifiable information was disclosed throughout the data processing, ensuring that our data usage is completely justifiable and legal.
@@ Appendix C Fine-Tuning Evaluation
Appendix D Additional Ablation Studies
D.1 Different LLMs for Attribute Construction.
To transform patient reports into structured descriptions of fundamental disease attributes, we employ a semi-automatic pipeline that uses GPT-4 (Achiam et al., 2023) for extracting lesion-related descriptions from each report, followed by further refinement by experienced human annotators. Given that using GPT-4 (Achiam et al., 2023) for attribute construction from radiological reports is often impractical due to privacy and legal concerns, we conduct extensive experiments to determine whether switching to locally hosted open-source LLMs can achieve comparable performance. The results are shown in Table 8. We compare the segmentation performance on unseen lesion categories using either the open-source medical large language model MMed-Llama-3-8B (Qiu et al., 2024) or GPT-4 (Achiam et al., 2023) to extract attribute descriptions from radiological reports.
Table 8: Ablation study on using different LLMs for attribute construction. As described in Sec. B.2, the LLMs are used only for our in-house dataset, as public datasets lack diagnostic reports and instead rely on attributes directly annotated by human experts.
Method In-house Dataset
Liver Cyst Kidney Stone Gallbladder Tumor
[…]
MMed-Llama-3-8B 54.97 63.06 38.66 47.71 43.28 51.45
GPT4 54.98 63.08 38.67 47.73 43.28 51.46
MMed-Llama-3-8B + Human Annotators 61.84 70.93 43.03 52.94 47.34 55.79
GPT4 + Human Annotators 61.85 70.93 43.05 52.95 47.35 55.79
The results indicate that using GPT-4 (Achiam et al., 2023) or MMed-Llama-3-8B (Qiu et al., 2024) for attribute extraction from radiological reports, with or without human annotation, has a negligible effect on model performance. This is largely because extracting and structuring attributes already present in the reports is a straightforward and easy task. However, incorporating human annotators significantly enhances performance, as some attributes are not explicitly mentioned in the reports and require expert supplementation—a task beyond the capabilities of current large language models. Given the high cost of human annotation, and to support future research, we will open-source our attribute annotations for all public datasets used in this study, along with our code, to benefit the research community.
D.2 The Number of Mask Tokens
An important hyperparameter in our framework is the number of mask tokens. We conducted ablation experiments to assess the segmentation performance of unseen lesions using different numbers of mask tokens on the KiTS23 dataset (Heller et al., 2023) and our in-house dataset. The results are presented in Fig. 4. We observed that when $N$ is less than 12, the model’s performance drops sharply. This observation is consistent with previous findings from MaskFormer-based methods (Cheng et al., 2021; 2022; Jiang et al., 2024), which suggest that the number of queries should be greater than the number of possible or useful classes in the data. When $N$ exceeds 16, no significant performance improvement was observed. Therefore, we set the default number of mask tokens to $N=16$.
Figure 4: Ablation studies on segmentation performance of unseen lesions using different numbers of mask tokens.
D.3 The Number of Attribute Aspects
This section addresses the key question of whether adding more attribute aspects improves zero-shot segmentation performance. We conducted ablation experiments to evaluate the segmentation performance of unseen lesions as we gradually added attribute categories in the order defined in Table 6. The results are presented in Fig. 5. We observed that as the number of attribute categories used for vision-language alignment increased from 1 to 8, the model’s zero-shot performance consistently improved with greater attribute diversity. This result demonstrates the superiority of our approach in using multiple attribute aspects to describe lesions from different perspectives, providing valuable textual features that help capture subtle visual characteristics of the lesions. This finding also suggests that incorporating additional contextual information or advanced features enhance the model’s ability to capture complex visual semantics.
Figure 5: Ablation studies on segmentation performance of unseen lesions using different numbers of attribute aspects.
Appendix E Evaluation of Lesion-Attribute Matching
@@
Figure 6: Visualization of segmentation and lesion-attribute matching results of Malenia. The numbers indicate the normalized similarity scores between the lesion masks and their corresponding attributes.
@@
Figure 7: Visualization illustrating how utilizing both mask and text embeddings improves segmentation performance for unseen and seen lesions with ambiguous boundaries. We present several prediction cases generated using both text embeddings and mask tokens, text embeddings only (TE), and mask tokens only (MT).
Appendix F Analysis of Model Robustness
F.1 Handling Lesions with Ambiguous Boundaries
Fig. 7 provides detailed examples illustrating how the CMKI module integrates text tokens and mask tokens to enhance the performance for lesions with ambiguous boundaries. We observed that when lesion boundaries are particularly blurry (as seen in the Hepatic Vessel Tumor cases on the left side of Fig. 7), the visual features from the mask tokens fail to accurately capture the lesion’s area. In such instances, text descriptions of attributes—such as the lesion’s shape and surface characteristics—provide additional information. The segmentation generated by the corresponding text embeddings refines the results produced by mask tokens alone, leading to improved performance.
Additionally, when encountering lesions with highly atypical appearances, such as colon tumors, which exhibit significant variability in shape and size as well as poorly defined boundaries (as seen in the colon tumor cases on the right side of Fig. 7), we observed that the model tends to produce more false positives when relying solely on mask tokens for segmentation predictions. Due to the clear semantic information and specific context provided by distinct attribute aspects, such as density, density variations, and surface texture, segmentation results generated using attribute embeddings exhibit significantly fewer false positives. These examples demonstrate how our framework leverages the strengths of text embeddings and mask tokens to address ambiguous or borderline cases, such as lesions with poorly defined boundaries or highly atypical appearances.
F.2 Handling Incorrect Text Inputs
During inference, we use all stored attribute embeddings. However, users may wish to specify their own attributes during testing, which could lead to the inclusion of incorrect attributes. In this section, we provide an example to illustrate how our model handles ambiguous or incorrect text inputs in Fig. 8. Specifically, we present two scenarios. (1) The model has access to both correct and incorrect text inputs (Fig. 8a). In this case, Malenia continues to produce stable segmentation results. Owing to our proposed mask-attribute alignment strategy, Malenia accurately computes the similarity between different text input embeddings and the mask tokens, effectively filtering out incorrect text inputs and selecting the best-matched attributes. (2) The model only has access to incorrect text inputs from users (Fig. 8b). For example, the image may contain a lung tumor in the right lung, but the user provides an attribute for the left lung. In this case, we visualized both the segmentation heatmap generated by the text embeddings and the one generated by the mask tokens. As shown in (Fig. 8b), although the text inputs are incorrect, the similarity between these inputs and the image features remains very low, resulting in segmentation heatmaps with low confidence. In contrast, the mask tokens continue to produce accurate segmentation heatmaps with high confidence. Thus, during the ensembling process, the high-confidence heatmaps from mask tokens prevent Malenia from being influenced by incorrect text inputs.
Appendix G Computational Cost Comparison
@@
Figure 8: Visualization of how Malenia handles incorrect text inputs.
Appendix H Discussion on Differences Between ZePT and Malenia
The primary distinctions between our approach and ZePT (Jiang et al., 2024) are as follows:
  • Alignment Strategy. ZePT Jiang et al. (2024) uses a single-scale mask-text alignment, aligning only the image features from the final Transformer decoder block with a basic textual description. In contrast, our method employs a multi-scale mask-attribute alignment, aligning multi-scale image features with detailed lesion attribute descriptions through a multi-positive contrastive loss. This represents a key technical advancement of our approach. The alignment strategy used by ZePT misses the opportunity to leverage multi-scale features, which hinders its ability to capture targets with significant scale variations. Furthermore, ZePT treats each mask and its corresponding category as the only positive sample pair, resulting in weak associations between different categories within the learned feature space. For instance, in ZePT, the mask corresponding to a liver tumor and the text embeddings of a liver cyst are treated as negative sample pairs, despite the shared attributes between liver tumors and liver cysts. Such similarities are overlooked by ZePT, whereas in our approach, these similarities are explicitly captured through attribute alignment. Therefore, Malenia demonstrates stronger generalization capabilities.
  • Mask Prediction. ZePT Jiang et al. (2024) relies solely on mask tokens for segmentation mask prediction. In contrast, our approach introduces a Cross-Modal Knowledge Injection (CMKI) module, which integrates features from both vision and language modalities, utilizing both text tokens and mask tokens for segmentation mask prediction. We demonstrate in Fig. 7 that CMKI is particularly effective in zero-shot lesion segmentation tasks, where unseen categories can be recognized by leveraging the complementary strengths of both vision and language modalities.
Thus, the main technical contributions of our work are: (1) Leveraging multi-scale features for cross-modal alignment. (2) Decomposing textual descriptions/reports into categorized attributes. (3) Introducing a multi-positive alignment mechanism that establishes better cross-modal feature representations. (4) Developing the Cross-Modal Knowledge Injection module. Our ablation studies also demonstrate that incorporating these designs significantly enhances zero-shot generalization performance. Furthermore, Fig. 7 shows that leveraging fine-grained textual information facilitates tumor segmentation, especially in scenarios where visual cues are limited.
Table 11: Details of the clinical knowledge table used during inference. We present several common diseases along with their corresponding attributes. We will open-source all diseases involved in this study along with their corresponding attributes.
Lesion Type Location Shape Density Relationship with Surrounding Organs
Hepatic Vessel Tumor Liver Round-like, Irregular Hypodense fluid-like lesion& Isodense lesion& Hyperdense lesion No close relationship with surrounding organs, Close relationship with adjacent organs
Pancreas Cyst Pancreas Cystic Hypodense fluid-like lesion No close relationship with surrounding organs
Kidney Tumor Kidney Round-like & Irregular Hypodense fluid-like lesion & Isodense lesion & Hyperdense lesion No close relationship with surrounding organs & Close relationship with adjacent organs
Liver Cyst Liver Cystic Hypodense fluid-like lesion No close relationship with surrounding organs
Kidney Stone Kidney Nodular Hyperdense lesion No close relationship with surrounding organs
Gallbladder Tumor Gallbladder Round-like & Irregular Hypodense lesion & Hyperdense lesion & Isodense lesion No close relationship with surrounding organs & Close relationship with adjacent organs
Lesion Type Density Variation Enhancement Status Surface Characteristics Specific Features
Hepatic Vessel Tumor Homogeneous & Heterogeneous Enhanced CT & Non-contrast CT Well-defined margin & Ill-defined margin Presence of decreased density areas & Presence of increased density areas
Pancreas Cyst Homogeneous Enhanced CT& Non-contrast CT Well-defined margin Cyst
Kidney Tumor Homogeneous & Heterogeneous Enhanced CT & Non-contrast CT Well-defined margin & Ill-defined margin Presence of decreased density areas & Presence of increased density areas
Liver Cyst Homogeneous Enhanced CT & Non-contrast CT Well-defined margin Cyst
Kidney Stone Homogeneous Enhanced CT & Non-contrast CT Well-defined margin Stone
Gallbladder Tumor Homogeneous & heterogeneous Enhanced CT & Non-contrast CT Well-defined margin & Ill-defined margin Stone
Appendix I Limitations and Future Work
Malenia currently demonstrates superior zero-shot lesion segmentation performance, nearing the performance of fully supervised, task-specific models. However, several limitations persist, revealing valuable directions for future research:
@@ Appendix I Limitations and Future Work
  • Cross-domain and cross-modality challenges: Zero-shot lesion segmentation across anatomical regions and imaging modalities remains highly challenging. For instance, training the model on lesions in abdominal CT data and directly testing it on brain cancer MRI is challenging due to significant differences in anatomical structures and feature distributions between the training and testing images. Achieving robust cross-domain and cross-modality zero-shot generalization is a complex but essential research direction in both natural and medical imaging domains.
  • Scalability to other imaging modalities: This work focuses primarily on 3D CT scans. While radiology reports for other modalities, such as X-ray and MRI, also contain structured attributes that could be leveraged, Malenia’s applicability to these modalities has not yet been tested. Future work will focus on adapting the framework to handle these imaging modalities, addressing their unique characteristics such as spatial resolution and semantic differences.
  • Handling lesions with ambiguous boundaries: Defining the boundaries of lesions with unclear or diffuse margins, such as pancreatic cancer or colon cancer, is inherently challenging and subjective even for radiologists. Our analysis in Table 9 indicates that Malenia’s mask-attribute matching performance on attributes such as Surface Characteristics still has room for improvement, as the inherent uncertainty in the corresponding ground truth labels affects segmentation accuracy. Future efforts will focus on enhancing the model’s ability to handle these ambiguous cases, potentially incorporating additional contextual information.
  • Capturing complex visual semantics: Malenia performs less effectively on attributes requiring intricate visual semantic understanding, such as Surface Characteristics and Specific Features. The challenges here may arise from the variability and imbalance in specific feature distributions, as well as the inherent difficulty in categorizing rare or atypical features. Exploring advanced model architectures and integrating additional contextual information will be crucial for enhancing performance on these attributes.
  • Generalization to diverse lesion types: While Malenia demonstrates strong performance on 12 lesion categories, the diversity of lesions in clinical practice is vast. Future work will focus on evaluating Malenia on a broader range of lesion types to further validate its generalization capabilities and ensure applicability to diverse clinical scenarios.
We believe addressing these limitations will significantly enhance Malenia’s robustness, scalability, and clinical utility, contributing to the broader advancement of zero-shot learning in medical imaging.